A planet of mass m = 1.55 x 1024 kg is orbiting in a circular path a star of mass M = 9.75 x 1029 kg. The radius of the orbit is R = 4.65 x 107 km. What is the orbital period (in earth days) of the planet? Where G = 6.67 � 10-11 N�m2/kg2 and 1 day = 8.54 � 104 s.

I used the formula
\frac{T^2}{r^3} = \frac{4\pi^2}{GM} and got T = 247 s = 0.0289 days, but the online homework system says thats the wrong answer.

What am I doing wrong?
Answer #1

T = 2*pi sqrt(r^3 / GM)
R=4.65*10^7km= 4.65*10^10metre

M=mass of star=9.75 x 1029 kg

T=7686902.828 sec or 90.01 days make sure u converted the radius in metres and took mass of star
Answer #2

The gravitational potential energy is equal to the centripetal force GmM mi 7 7 GM GM 4T273 GM GM Substitute the given value
Answer #3

T = 2*pi sqrt(r^3 / GM)
R=4.65*10^7km= 4.65*10^10metre

M=mass of star=9.75 x 1029 kg

G= 6.67